OpenStudy (anonymous):

The set of x-intercepts of the graph of f(x) = x^3 -2x^2 -x +2 is?

OpenStudy (anonymous):

set f(x) to 0. 0=x^3-2x^2-x+2 by factoring out the right hand side, you get 0= (-2+x)(-1+x)(1+x), solve for x which matches the solution of 0.

OpenStudy (anonymous):

Could you show me with steps how you factored the right hand side? thats what i had trouble with

OpenStudy (loser66):

use synthetic method.

OpenStudy (anonymous):

Can you show me that?

OpenStudy (loser66):

read it, at the middle of the page. The first part is about long division, skip it.

OpenStudy (loser66):

http://www.purplemath.com/modules/synthdiv.htm

hero (hero):

f(x) = x^3 -2x^2 -x +2 0 = x^3 - 2x^2 - x + 2 0 = x^2(x - 2) - 1(x - 2) 0 = (x - 2)(x^2 - 1) 0 = (x - 2) 0 = (x^2 - 1) `2` = x 1 = x^2 `±1` = x